OpenStudy (anonymous):

We roll a pair of dice. If the sum of the dice is 7, you pay me $26. If the sum is not 7, I pay you the number of dollars indicated by the sum of the dice. What is your expected value for the game?

OpenStudy (anonymous):

we got to multiply the amount you win by the probability you win it and add it all up

OpenStudy (anonymous):

wait so 26 * 7?

OpenStudy (anonymous):

no it has nothing to do with \(26\times 7\)

OpenStudy (anonymous):

ok im confused

OpenStudy (anonymous):

yes, i see that

OpenStudy (anonymous):

do you know the probability you roll a 7 on two dice?

OpenStudy (anonymous):

if not, we can start there, because you are going to need this you are going to need the probability for every possible roll of the dice actually do you know them? "no" is a fine answer, we can find them all

OpenStudy (anonymous):

i think so but im not positive. all i remember was there was a chart that listed the probabilities or at least the amount when rolling two die....

OpenStudy (anonymous):

yeah you got that chart? we will need it

OpenStudy (anonymous):

yes i can look it up

OpenStudy (anonymous):

ok good all the denominators we will leave as \(36\) because you are going to have to add

OpenStudy (anonymous):

ok that makes sense.

OpenStudy (anonymous):

btw thank you for being understanding, all the other tutors have made me feel so dumb and i honestly just need help cause ive taken this class like twice already and im beyond brain fried on the subject

OpenStudy (anonymous):

also, you do understand that if you roll a 2 you get $2, if you roll a 3 you get $3, if you roll a 4 you get $4 etc, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so our job it so multiply the amount you win by the probability you win it, and add up all the numbers

OpenStudy (anonymous):

i happen to remember them the probability you roll a 2 is \(\frac{1}{36}\) so one term will be \(2\times \frac{1}{36}=\frac{2}{36}\) but lets just leave it as \(2\times \frac{1}{36}\)

OpenStudy (anonymous):

the next one will be \(3\times \frac{2}{36}\) because the probability you roll a 3 is \(\frac{2}{36}\) you following along?

OpenStudy (anonymous):

yeah so then would you add: 1/36 + 2/36 +3/36 etc...

OpenStudy (anonymous):

careful lets go slow

OpenStudy (anonymous):

it is not what you wrote, it is this \[2\times \frac{1}{36}+3\times \frac{2}{36}+4\times \frac{3}{36}+5\times \frac{4}{36}+6\times \frac{5}{36}\] there is more

OpenStudy (anonymous):

oh!! ok

OpenStudy (anonymous):

\[\overbrace{2}^{\text{ amount you win}}\times \overbrace{\frac{1}{36}}^{\text{probability you win it}}+3\times \frac{2}{36}+4\times \frac{3}{36}+5\times \frac{4}{36}+6\times \frac{5}{36}\]

OpenStudy (anonymous):

all the way down the line lets leave off the rolling 7 for a moment, you lose for that one

OpenStudy (anonymous):

so we have more terms \[8\times \frac{5}{36}+9\times \frac{4}{36}+10\times \frac{3}{36}+11\times \frac{2}{36}+12\times \frac{1}{36}\]

OpenStudy (anonymous):

the denomnators are all 36 so we just need to add up in the numerator \[\frac{2\times 1+3\times 2+4\times 3+5\times 4+6\times 5+8\times 5+9\times 4+10\times 3+11\times 2+12\times 1}{36}\]

OpenStudy (anonymous):

i told you this was a pain but not terribly hard just annoying lets use a calculator

OpenStudy (anonymous):

i get \(210\) http://www.wolframalpha.com/input/?i=2%2B3*2%2B4*3%2B5*4%2B6*5%2B8*5%2B9*4%2B10*3%2B11*2%2B12 so \[\frac{210}{36}\] then we have to subtract the amount we lose times the probability we lose it

OpenStudy (anonymous):

\[\frac{210}{36}-26\times \frac{6}{36}\]

OpenStudy (anonymous):

\[\frac{210-156}{36}\] is our answer

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

make sure to pass this time!

OpenStudy (anonymous):

hahah no kidding!!